# Graph Anomaly Detection via Adaptive Test-time Representation Learning across Out-of-Distribution Domains

### v1

@@ Abstract
Graph Anomaly Detection (GAD) has demonstrated great effectiveness in identifying unusual patterns within graph-structured data. However, while labeled anomalies are often scarce in emerging applications, existing supervised GAD approaches are either ineffective or not applicable when moved across graph domains due to distribution shifts and heterogeneous feature spaces. To address these challenges, we present AdaGraph-T3, a novel test-time training framework for cross-domain GAD. AdaGraph-T3 combines supervised and self-supervised learning during training while adapting to a new domain during test time using only self-supervised learning by leveraging a homophily-based affinity score that captures domain-invariant properties of anomalies. Our framework introduces four key innovations to cross-domain GAD: an effective self-supervision scheme, an attention-based mechanism that dynamically learns edge importance weights during message passing, domain-specific encoders for handling heterogeneous features, and class-aware regularization to address imbalance. Experiments across multiple cross-domain settings demonstrate that AdaGraph-T3 significantly outperforms existing approaches, achieving average improvements of over 6.6% in AUROC and 7.9% in AUPRC compared to the best competing model.
@@ 1 Introduction
Graph Anomaly Detection (GAD) is a critical task to identify unusual patterns (or outliers) in graph-structured data [24, 2]. This problem has many real-world applications such as in e-commerce [45], social networks [33], fraud detection [12], and cybersecurity [17]. A key limitation of existing GAD models is that they face unique cross-domain challenges that distinguish them from general graph learning tasks. These include inconsistent definitions of normal/anomalous patterns and heterogeneous feature spaces between domains (e.g., Facebook vs. Amazon datasets). These applications would benefit from GAD models capable of adapting across Out-Of-Distribution (OOD) and heterogeneous graphs.
@@
A major motivation for our work is cybersecurity. Network intrusion detection systems play a key role in identifying malicious network activity associated with cyberattacks [31, 14]. Due to the volume and dynamic nature of these attacks, modern intrusion detection systems increasingly rely on large amounts of data and machine learning to assist cybersecurity experts in detecting potential intrusions. However, one of the major challenges in data-driven cybersecurity is the lack of sufficient labeled data for training supervised models. This is particularly critical for emerging applications, where traces of labeled intrusions are scarce and the ability to leverage attacks or other types of anomalous behavior from existing labeled datasets would be greatly beneficial.
In this paper, we investigate the task of unsupervised domain adaptation for GAD focusing on the transfer of knowledge from a labeled source to an unlabeled target domain. While domain adaptation for graphs has been extensively studied [29, 37, 47, 42, 35, 9], cross-domain graph anomaly detection is an emerging area [36, 8]. We propose using Test-Time Training (TTT) [30] to address key challenges in cross-domain graph anomaly detection. Unlike traditional domain adaptation methods that require continuous access to source data, TTT allows us to distill source knowledge into model parameters and adapt to target domains using only self-supervision. This is particularly advantageous for graph anomaly detection where source data may be sensitive or unavailable during the adaptation phase to new domains.
@@
Our approach, Adaptive Graph Test-Time Training (AdaGraph-T3), has four major innovations compared to existing work on cross-domain GAD. First, it leverages a homophily-based affinity score [26, 4] for self-supervised learning based on the observation that normal nodes are more similar to their neighbors than anomalous ones, a pattern that remains consistent across domains (see Figure 1). Second, it introduces Normal Structure-preserved Attention Weighting (NSAW), which dynamically learns continuous edge importance weights through the attention mechanism. Third, AdaGraph-T3 applies source and target-specific encoders that are trained end-to-end to handle both distribution shifts and heterogeneous feature spaces. Fourth, to address the extreme class imbalance challenge in GAD [23], our approach employs class-aware regularization during source training—with a stronger regularization to the minority class.
@@
We compare our solution against both graph domain adaptation and graph anomaly detection approaches from the literature using multiple cross-domain datasets. For instance, we show that labeled anomalies in the Amazon dataset (source) can improve the GDA accuracy on the Facebook dataset (target). The experiments show that AdaGraph-T3 significantly outperforms the alternatives in most of the settings. The contributions of this paper can be summarized as:
@@
  • We propose the first test-time training (TTT) framework for cross-domain graph anomaly detection (AdaGraph-T3). TTT enables adaptation to new target domains without requiring direct access to source data or target labels at test time. Our approach leverages homophily-based self-supervision which captures universal patterns, as anomalous nodes typically exhibit lower homophily compared to normal nodes.
@@
  • We propose several strategies specifically designed for the cross-domain GAD problem such as data-specific encoders, class-aware regularization, and an effective early-stopping strategy to prevent overfitting during test-time adaptation.
  • AdaGraph-T3’s performance is demonstrated through extensive experiments on cross-network tasks using multiple datasets and baselines that span both the domain adaptation and the anomaly detection literature. Our approach achieves average improvements of over 6.6% in AUROC and 7.9% in AUPRC compared to the best competing model.
@@ 2.1 Graph Domain Adaptation (GDA)
While graph domain adaptation (GDA) methods have shown success in transferring knowledge across different domains [29], they face fundamental limitations when applied to anomaly detection tasks. As traditional GDA methods are designed for node classification scenarios with balanced classes and consistent label semantics.
Traditional Graph Domain Adaptation (GDA) approaches, such as GRADE [39], AdaGCN [6], and UDA-GCN [40] and spectral regularization methods [44] focus on minimizing distribution shifts across domains but require training samples during inference, limiting their real-time adaptation capabilities. Test-time training (TTT) [30] has emerged as a powerful framework for handling distribution shifts [19, 41, 46]. Methods like GTrans [13], GraphCL [43], and GT3 [38] propose graph-specific approaches. TENT [34] uses prediction entropy minimization, while GraphTTA [3] leverages information theory for TTT on graphs. However, existing GDA approaches, including both traditional methods and TTT-based solutions, face key limitations for anomaly detection as they (1) don’t account for extreme class imbalance within domains, (2) lack mechanisms to preserve anomaly-indicating patterns during transfer, and (3) don’t handle heterogeneous feature spaces across domains. While there exists limited research on cross-domain graph anomaly detection, prior approaches require direct access to the source dataset during adaptation [8, 36]. Our approach offers key advantages over these methods by eliminating the need to store and access the full source dataset during adaptation. This makes AdaGraph-T3 both privacy-preserving and memory-efficient, making it more practical for real-world deployment where data privacy and computational resources are concerns.
Our work addresses key challenges in deploying GDA in anomaly detection through dataset-specific encoders for heterogeneous feature spaces and a homophily-based unsupervised learning approach that extracts domain-invariant properties of anomalies that generalize across different graph domains. In our experiments, we show that AdaGraph-T3 often outperforms multiple GDA baselines, including GRADE, AdaGCN, TENT, and GTrans.
Figure 1: Distributions of homophily scores show normal nodes (blue) higher than anomalous nodes (orange) across domains, indicating homophily as a domain-invariant anomaly detector.
@@ 2.2 Graph Anomaly Detection (GAD)
Graph Anomaly Detection (GAD) [24, 2] focuses on identifying abnormal nodes in graph-structured data. Traditional approaches like Oddball [1] rely on power-law relationships between local graph features, while more recent deep learning-based approaches are more generalizable. For instance, DOMINANT [7] employs a graph autoencoder to identify anomalies based on graph reconstruction. ComGA [22] introduces a tailored GCN to learn distinguishable node representations by explicitly capturing community structure. Self-supervised techniques have emerged as powerful tools for GAD, with methods like CoLA [21], SL-GAD [48], HCM-A [11], and TAM [26] introducing various approaches to handle node interactions and structural patterns. While these methods have shown success in single-domain scenarios, they don’t address the challenges of cross-domain knowledge transfer. First, anomalies behave differently across domains– fraudulent users in e-commerce networks exhibit different patterns compared to social networks. Second, domains often have different feature spaces and graph structures.
AdaGraph-T3 bridges both GDA and GAD through a unified approach that combines (1) test-time training with a homophily-based loss for capturing domain-invariant properties of anomalies (2) enhanced message-passing using normal structure-preserved attention weighting (NSAW) that reduces the irrelevant impact of anomalous nodes on normal nodes’ representations, and (3) class-aware regularization that prevents minority class patterns from being overshadowed by majority class during source training. We compare a source-free version of AdaGraph-T3 against GAD baselines (e.g. TAM, ComGA, DOMINANT) to highlight the impact of its GAD-specific features.
@@ 3 Problem Definition
We address unsupervised node-level anomaly detection with domain adaptation, aiming to identify abnormal nodes in a target graph by leveraging information from both the target and a source graph. The source dataset may contain labeled information but is assumed to be out-of-distribution (OOD) relative to the target, with different feature sets and label distributions. Given source dataset $D_{s}=(\mathcal{G}_{s},X_{s},Y_{s})$ and target dataset $D_{t}=(\mathcal{G}_{t},X_{t})$, where $\mathcal{G}_{s}/\mathcal{G}_{t}=(\mathcal{V},\mathcal{E})$ represents a (source or target) underlying graph with nodes $\mathcal{V}$ and edges $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ such that $uv\in\mathcal{E}$ if there is a link between nodes $u$ and $v$. Here, $X_{s}=\{\mathbf{x}_{v}|\,\forall v\in\mathcal{V}_{s}\}$ and $Y_{s}=\{y_{v}|\,\forall v\in\mathcal{V}_{s}\}$ represent the node features and their corresponding labels, respectively, and the same holds for target data $(X_{t},Y_{t})$. Notably, $y_{v}\in\{0,1\}$ (normal or anomaly) for nodes in both $\mathcal{V}_{s}$ and $\mathcal{V}_{t}$. Our goal is to detect anomalies in the target graph. The key challenge is that target labels are completely unavailable (unsupervised setting), requiring us to leverage source labels $Y_{s}$ as auxiliary information, even though source and target features may have different dimensionalities ($\mathbf{x}_{v}\in\mathbb{R}^{p_{s}}$ for source nodes and $\mathbf{x}_{u}\in\mathbb{R}^{p_{t}}$ for target nodes).
4 Adaptive Graph Anomaly Detection via Test-Time Training (AdaGraph-T3)
@@
We propose AdaGraph-T3 (Adaptive Graph Anomaly Detection via Test-Time Training), a domain adaptation method particularly designed for graph anomaly detection that integrates attention-based message-passing, node homophily patterns, and domain-specific encoders to handle feature shifts and heterogeneous features across domains. AdaGraph-T3 addresses the challenging and novel scenario where the target data is out of distribution (OOD) relative to the source data and where source and target datasets have different feature spaces.
@@
An overview of our approach is provided in Figure 2. AdaGraph-T3 leverages the advantages of Test-time Training, especially the ability to adapt to new target datasets in the TTT phase without the need for labeled anomalies. Moreover, AdaGraph-T3 incorporates technical innovations tailored for cross-domain GAD, including the use of domain-specific encoders, Normal Structure-preserved Attention Weighting (NSAW), and class-aware regularization.
@@
Our framework operates in two phases. The training phase combines supervised and self-supervised losses to learn from the source dataset. The test-time training phase only uses self-supervised learning to adapt to the target data. We will detail each stage of our solution together with our key contributions in the next sections.
@@
Figure 2: Overview of AdaGraph-T3, our test-time training framework for node-level cross-domain anomaly detection. In the training phase (top), AdaGraph-T3 learns an encoder $\theta_{s}$, a decoder $\theta_{m}$, and a prediction $\theta_{pred}$ by jointly minimizing supervised ($\mathcal{L}_{\text{sup}}$) and self-supervised homophily-based loss ($\mathcal{L}_{\text{self}}$). We propose using Normal Structure-preserved Attention Weights to defend against the adversarial influence of anomalous nodes during message passing in the GNN training. In the test-time training (TTT) phase (bottom), AdaGraph-T3 learns a target encoder $\theta_{t}$ using only the self-supervised loss $\mathcal{L}_{\text{self}}$. The decoder $\theta_{m}$ and prediction $\theta_{pred}$ are shared across phases to enable the cross-domain knowledge transfer while handling heterogeneous feature spaces.

### v2

@@ Abstract
Graph Anomaly Detection (GAD) has demonstrated great effectiveness in identifying unusual patterns within graph-structured data. However, while labeled anomalies are often scarce in emerging applications, existing supervised GAD approaches are either ineffective or not applicable when moved across graph domains due to distribution shifts and heterogeneous feature spaces. To address these challenges, we present GADT3, a novel test-time training framework for cross-domain GAD. GADT3 combines supervised and self-supervised learning during training while adapting to a new domain during test time using only self-supervised learning by leveraging a homophily-based affinity score that captures domain-invariant properties of anomalies. Our framework introduces four key innovations to cross-domain GAD: an effective self-supervision scheme, an attention-based mechanism that dynamically learns edge importance weights during message passing, domain-specific encoders for handling heterogeneous features, and class-aware regularization to address imbalance. Experiments across multiple cross-domain settings demonstrate that GADT3 significantly outperforms existing approaches, achieving average improvements of over 8.2% in AUROC and AUPRC compared to the best competing model. The source code for GADT3 is available at https://github.com/delaramphf/GADT3-Algorithm.
@@ 1 Introduction
Graph Anomaly Detection (GAD) is a critical task to identify unusual patterns (or outliers) in graph-structured data (Ma et al., 2021; Akoglu et al., 2015). This problem has many real-world applications, such as in e-commerce (Zhang et al., 2022), social networks (Venkatesan & Prabhavathy, 2019), fraud detection (Jiang et al., 2019), and cybersecurity (Lazarevic et al., 2003). A key limitation of existing GAD models is that they face unique cross-domain challenges that distinguish them from general graph learning tasks. These include inconsistent definitions of normal/anomalous patterns and heterogeneous feature spaces between domains (e.g., Facebook vs. Amazon datasets). These applications would benefit from GAD models capable of adapting across Out-Of-Distribution (OOD) and heterogeneous graphs.
@@
A major motivation for our work is cybersecurity. Network intrusion detection systems play a key role in identifying malicious network activity associated with cyberattacks (Tsai et al., 2009; Kilincer et al., 2021). Due to the volume and dynamic nature of these attacks, modern intrusion detection systems increasingly rely on large amounts of data and machine learning to assist cybersecurity experts in detecting potential intrusions. However, one of the major challenges in data-driven cybersecurity is the lack of sufficient labeled data for training supervised models. This is particularly critical for emerging applications, where traces of labeled intrusions are scarce, and the ability to leverage attacks or other types of anomalous behavior from existing labeled datasets would be greatly beneficial.
@@
This paper investigates the unsupervised domain adaptation problem for GAD, focusing on the transfer of knowledge from a labeled source to an unlabeled target domain. While domain adaptation for graphs has been studied by previous work (Shi et al., 2024; Wang et al., 2021; Zheng et al., 2019; Yin et al., 2023; Wang et al., 2023a; Ding et al., 2018), cross-domain graph anomaly detection is an emerging challenge (Wang et al., 2023b; Ding et al., 2021). We propose using Test-Time Training (TTT) (Sun et al., 2020) to address cross-domain graph anomaly detection. Unlike traditional domain adaptation methods that require continuous access to source data, TTT allows us to distill source knowledge into model parameters and adapt to target domains using only self-supervision. This is particularly advantageous for graph anomaly detection, where source data may be sensitive or unavailable during the adaptation phase. Our approach, Graph Anomaly Detection with Test-time Training (GADT3), has four major innovations compared to existing work on cross-domain GAD. First, it leverages a homophily-based affinity score (Qiao & Pang, 2024; Chen et al., 2024b) for self-supervised learning based on the observation that normal nodes are more similar to their neighbors than anomalous ones, a pattern that remains consistent across domains. We observe a clear separation in homophily scores between normal and anomalous nodes across most datasets, justifying the use of homophily-based SSL. Figure 1 shows the homophily score distributions for two datasets (see Appendix A.11 for detailed plots). Second, it introduces Normal Structure-preserved Attention Weighting (NSAW), which dynamically learns continuous edge importance weights through the attention mechanism to suppress the effect of anomalous nodes on the representation of normal ones. Third, GADT3 applies source and target-specific encoders that are trained end-to-end to handle both distribution shifts and heterogeneous feature spaces. Fourth, to address the extreme class imbalance challenge in GAD (Ma et al., 2024), our approach employs class-aware regularization during source training—with a stronger regularization to the minority class. By combining these innovations, GADT3 can effectively identify anomalous nodes across heterogeneous datasets arising from multiple and diverse domains.
@@
We compare our solution against both graph domain adaptation and graph anomaly detection approaches using multiple cross-domain datasets. For instance, we show that labeled anomalies in the Amazon dataset (source) can improve the GAD accuracy on the Facebook dataset (target). The experiments show that GADT3 significantly outperforms the alternatives in most of the settings. The contributions of this paper can be summarized as:
@@
  • We propose the first test-time training (TTT) framework for cross-domain graph anomaly detection (GADT3). TTT enables adaptation to new target domains without requiring direct access to source data or target labels at test time. Our approach leverages homophily-based self-supervision, which captures universal patterns, as anomalous nodes typically exhibit lower homophily compared to normal nodes.
@@
  • We propose several strategies specifically designed for cross-domain GAD, such as data-specific encoders, class-aware regularization, and an early-stopping strategy to prevent overfitting during test-time adaptation.
  • GADT3’s performance is demonstrated through extensive experiments on cross-network tasks using multiple datasets and baselines that span both domain adaptation and anomaly detection. Our approach achieves average relative improvements of over 8.2% in AUROC and AUPRC compared to the best competing model.
@@ 2.1 Graph Domain Adaptation (GDA)
While graph domain adaptation (GDA) methods have shown success in transferring knowledge across different domains (Shi et al., 2024), they face fundamental limitations when applied to anomaly detection tasks. Traditional GDA methods are designed for node classification scenarios with balanced classes and consistent label semantics.
Traditional Graph Domain Adaptation (GDA) approaches, such as GRADE (Wu et al., 2023), AdaGCN (Dai et al., 2022), and UDA-GCN (Wu et al., 2020), and spectral regularization methods (You et al., 2023), focus on minimizing distribution shifts across domains but require training samples during inference, limiting their real-time adaptation capabilities. Test-time training (TTT) (Sun et al., 2020) has emerged as a powerful framework for handling distribution shifts (Li et al., 2022; Wu et al., 2024; Zhang et al., 2024; Liu et al., 2021b). Methods like GTrans (Jin et al., 2022), GraphCL (You et al., 2020), and GT3 (Wang et al., 2022) propose graph-specific TTT approaches. TENT (Wang et al., 2020) uses prediction entropy minimization, while GraphTTA (Chen et al., 2022) leverages information theory for TTT on graphs. However, existing GDA approaches, including both traditional methods and TTT-based solutions, face key limitations for anomaly detection as they (1) don’t account for extreme class imbalance within domains, (2) lack mechanisms to preserve anomaly-indicating patterns during transfer, and (3) don’t handle heterogeneous feature spaces across domains. While there exists limited research on cross-domain graph anomaly detection, prior approaches require direct access to the source dataset during adaptation (Ding et al., 2021; Wang et al., 2023b). Our approach offers key advantages over these methods by eliminating the need to store and access the full source dataset during adaptation. This makes GADT3 both privacy-preserving and memory-efficient, thus being more practical for real-world deployment where these are major concerns.
Our work addresses key challenges in deploying GDA in anomaly detection through dataset-specific encoders for heterogeneous feature spaces and a homophily-based unsupervised learning approach that extracts domain-invariant properties of anomalies that generalize across domains. In our experiments, we show that GADT3 often outperforms multiple GDA baselines, including GRADE, AdaGCN, TENT, and GTrans.
Figure 1: Homophily score distributions across domains (Amazon to Reddit): normal nodes (blue) show higher scores than anomalous nodes (orange), suggesting homophily as a domain-invariant anomaly signal.
@@ 2.2 Graph Anomaly Detection (GAD)
Graph Anomaly Detection (GAD) (Ma et al., 2021; Akoglu et al., 2015) focuses on identifying abnormal nodes in graph-structured data. Traditional approaches like Oddball (Akoglu et al., 2010) rely on power-law relationships between local graph features, while more recent deep learning-based approaches are more generalizable. For instance, DOMINANT (Ding et al., 2019) employs a graph autoencoder to identify anomalies based on graph reconstruction. ComGA (Luo et al., 2022) introduces a tailored GCN to learn distinguishable node representations by explicitly capturing community structure. Self-supervised techniques have emerged as powerful tools for GAD, with methods like CoLA (Liu et al., 2021a), SL-GAD (Zheng et al., 2021), HCM-A (Huang et al., 2022), and TAM (Qiao & Pang, 2024) introducing various approaches to handle node interactions and structural patterns. While these methods have shown success in single-domain scenarios, they don’t address the challenges of cross-domain knowledge transfer. Recent methods like ACT (Wang et al., 2023b) and COMMANDER (Ding et al., 2021) have started to explore cross-domain graph anomaly detection (CD-GAD), but this remains an evolving field. Other recent works like CDFS-GAD (Chen et al., 2024a) address cross-domain node anomaly detection in a few-shot setting, where a small set of labeled anomalies from the target graph is available. ARC (Liu et al., 2024) proposes an in-context learning approach for cross-domain GAD, which adapts at inference using a few normal nodes from the target graph.
Anomalies behave differently across domains—fraudulent users in e-commerce networks exhibit different patterns compared to social networks. Second, domains often have different feature spaces and graph structures. While our method is motivated by the homophily assumption, recent studies such as (Gao et al., 2023) have explored anomaly detection in heterophilic graphs, highlighting the importance of developing models robust to varying structural patterns. Our approach can be extended to heterophilic graphs by replacing the current SSL loss with a loss more suited for heterophily, and by using heterophily-aware GNNs.
GADT3 bridges both GDA and GAD through a unified approach that combines (1) test-time training with a homophily-based loss for capturing domain-invariant properties of anomalies (2) enhanced message-passing using normal structure-preserved attention weighting (NSAW) that reduces the irrelevant impact of anomalous nodes on normal nodes’ representations, and (3) class-aware regularization that prevents minority class patterns from being overshadowed by the majority class during source training. In our experiments, we compare a source-free version of GADT3 against GAD baselines (e.g., TAM, ComGA, DOMINANT) to highlight the impact of its GAD-specific features.
@@ 3 Problem Definition
We address unsupervised node-level anomaly detection with domain adaptation, aiming to identify abnormal nodes in a target graph by leveraging information from both the target and a source graph. The source dataset contains labeled information but is assumed to be out-of-distribution (OOD) relative to the target, with different feature sets and label distributions. Given source dataset $D_{s}=(\mathcal{G}_{s},X_{s},Y_{s})$ and target dataset $D_{t}=(\mathcal{G}_{t},X_{t})$, where $\mathcal{G}_{s}/\mathcal{G}_{t}=(\mathcal{V},\mathcal{E})$ represents a (source or target) underlying graph with nodes $\mathcal{V}$ and edges $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ such that $uv\in\mathcal{E}$ if there is a link between nodes $u$ and $v$. Here, $X_{s}=\{\mathbf{x}_{v}|\,\forall v\in\mathcal{V}_{s}\}$ and $Y_{s}=\{y_{v}|\,\forall v\in\mathcal{V}_{s}\}$ represent the node features and their corresponding labels, respectively, and the same holds for target data $(X_{t},Y_{t})$. Notably, $y_{v}\in\{0,1\}$ (normal or anomaly) for nodes in both $\mathcal{V}_{s}$ and $\mathcal{V}_{t}$. Our goal is to detect anomalies in the target graph. The key challenge is that target labels are completely unavailable (unsupervised setting), requiring us to leverage source labels $Y_{s}$ as auxiliary information, even though source and target features may have different dimensionalities and semantic meanings ($\mathbf{x}_{v}\in\mathbb{R}^{p_{s}}$ for source nodes and $\mathbf{x}_{u}\in\mathbb{R}^{p_{t}}$ for target nodes).
4 Graph Anomaly Detection via Test-Time Training (GADT3)
@@
We propose GADT3 (Graph Anomaly Detection with Test-time Training), a domain adaptation method particularly designed for graph anomaly detection that integrates attention-based message-passing, node homophily patterns, and domain-specific encoders to handle feature shifts and heterogeneous features across domains. GADT3 addresses the challenging and novel scenario where the target data is out of distribution (OOD) relative to the source data and where source and target datasets have different feature spaces.
@@
An overview of our approach is provided in Figure 2. GADT3 leverages the advantages of Test-time Training, especially the ability to adapt to new target datasets in the TTT phase without the need for labeled anomalies. Moreover, GADT3 incorporates technical innovations tailored for cross-domain GAD, including the use of domain-specific encoders, Normal Structure-preserved Attention Weighting (NSAW), and class-aware regularization. Unlike prior TTT methods that rely on auxiliary dummy tasks (e.g., rotation prediction or jigsaw solving) for encoder adaptation, our self-supervised homophily loss is directly aligned with the target anomaly detection objective. By encouraging high homophily among node embeddings in the target graph, this task-relevant signal promotes local consistency, enabling the model to adapt without labels while preserving the semantic structure essential for anomaly detection.
@@
Our framework operates in two phases. The training phase combines supervised and self-supervised losses to learn from the source dataset. The test-time training phase only uses self-supervised learning to adapt to the target data. During training, we learn the encoder $\theta_{s}$, decoder $\theta_{m}$, and supervised predictor $\theta_{\text{pred}}$ jointly based on the source dataset. During test-time training (TTT) on the target domain, a new encoder $\theta_{t}$ is trained from scratch using the unsupervised homophily-based loss $\mathcal{L}_{\text{self}}$, while $\theta_{m}$ is reused from the training phase. This decoupled encoder-decoder structure enables adaptation without requiring any target labels or access to source data after training. We will detail each stage of our solution together with our key contributions in the next sections.
@@
Figure 2: Overview of GADT3, our test-time training framework for node-level cross-domain anomaly detection. In the training phase (top), GADT3 learns an encoder $\theta_{s}$, a decoder $\theta_{m}$, and a prediction $\theta_{pred}$ by jointly minimizing supervised ($\mathcal{L}_{\text{sup}}$) and self-supervised homophily-based loss ($\mathcal{L}_{\text{self}}$). We propose using Normal Structure-preserved Attention Weights to defend against the adversarial influence of anomalous nodes during message passing in the GNN training. In the test-time training (TTT) phase (bottom), GADT3 learns a target encoder $\theta_{t}$ using only the self-supervised loss $\mathcal{L}_{\text{self}}$. The decoder $\theta_{m}$ is shared across phases to enable the cross-domain knowledge transfer while handling heterogeneous feature spaces. During inference, the adapted target representations produced by $\theta_{m}$ are used for anomaly scoring.